### v3

@@ 4.2 Minimizing a random complex quartic polynomial with unit-norm variables
The second example is to minimize a random complex quartic polynomial with unit-norm variables:
@@
We approach (11) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 3. From the table, we see that the reformulation ($-{\mathbb{R}}$) is about one magnitude faster than the reformulation ($-{\mathbb{R}}$), and the speedup becomes more significant as the SDP size grows. Again, solving the original complex SDP with Hypatia is much slower than solving the real SDP reformulation ($-{\mathbb{R}}$) with MOSEK.
@@
$s$ $d$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
@@ 4.3 Minimizing a randomly generated sparse complex quartic polynomial over multi-spheres
The sparsity in (12) can be exploited to derive a sparsity-adapted complex moment-HSOS hierarchy [15]. We solve the second sparse HSOS relaxation of (12) for $l=40,80,\ldots,400$. The results are displayed in Table 4. From the table we see that the reformulation ($-{\mathbb{R}}$) is $1.5\sim 2$ times as fast as the reformulation ($-{\mathbb{R}}$). Moreover, for this problem, solving the original complex SDP with Hypatia is extremely slow probably due to the fact that the SDP contains many PSD blocks.
@@
$l$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$) ($-{\mathbb{C}}$)
@@ 4.4 Application to the AC-OPF problem
(13) $\begin{cases}\inf\limits_{V_{i},S_{k}^{g}}&\sum_{k\in G}\left(\mathbf{c}_{2k}({\mathcal{R}}(S_{k}^{g}))^{2}+\mathbf{c}_{1k}{\mathcal{R}}(S_{k}^{g})+\mathbf{c}_{0k}\right)\\ \,\,\,\textrm{s.t.}&\angle V_{r}=0,\\ &\mathbf{S}_{k}^{gl}\leq S_{k}^{g}\leq\mathbf{S}_{k}^{gu},\quad\forall k\in G,\\ &\boldsymbol{\upsilon}_{i}^{l}\leq|V_{i}|\leq\boldsymbol{\upsilon}_{i}^{u},\quad\forall i\in N,\\ &\sum_{k\in G_{i}}S_{k}^{g}-\mathbf{S}_{i}^{d}-\mathbf{Y}_{i}^{sh}|V_{i}|^{2}=\sum_{(i,j)\in E_{i}\cup E_{i}^{R}}S_{ij},\quad\forall i\in N,\\ &S_{ij}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})\frac{|V_{i}|^{2}}{|\mathbf{T}_{ij}|^{2}}-\overline{\mathbf{Y}}_{ij}\frac{V_{i}\overline{V}_{j}}{\mathbf{T}_{ij}},\quad\forall(i,j)\in E,\\ &S_{ji}=(\overline{\mathbf{Y}}_{ij}-\mathbf{i}\frac{\mathbf{b}_{ij}^{c}}{2})|V_{j}|^{2}-\overline{\mathbf{Y}}_{ij}\frac{\overline{V}_{i}V_{j}}{\overline{\mathbf{T}}_{ij}},\quad\forall(i,j)\in E,\\ &|S_{ij}|\leq\mathbf{s}_{ij}^{u},\quad\forall(i,j)\in E\cup E^{R},\\ &\boldsymbol{\theta}_{ij}^{\Delta l}\leq\angle(V_{i}\overline{V}_{j})\leq\boldsymbol{\theta}_{ij}^{\Delta u},\quad\forall(i,j)\in E,\\ \end{cases}$
@@
We select test cases from the AC-OPF library PGLiB-OPF [3]. For each case, we solve the minimal relaxation step of the sparse HSOS hierarchy [15]. The results are displayed in Table 5. From the table, we again see that the reformulation ($-{\mathbb{R}}$) is several ($1.4\sim 5$) times as fast as the reformulation ($-{\mathbb{R}}$).
@@
Case $s$ $t$ ($-{\mathbb{R}}$) ($-{\mathbb{R}}$)
@@
[…]

### v4

@@ 4.3 Minimizing a random complex quartic polynomial with unit-norm variables
The next example is to minimize a random complex quartic polynomial with unit-norm variables:
@@
We approach (13) for $s=5,7,\ldots,15$ with the second and third HSOS relaxations. The related results are shown in Table 4. From the table, we see that the reformulation (HSOS-$\mathbb{R}$}) is about one magnitude faster than the reformulation (HSOS-$\mathbb{R}$}), and again the speedup becomes more significant as the SDP size grows.
@@
$s$ $d$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})
@@ 4.4 Minimizing a randomly generated sparse complex quartic polynomial over multi-spheres
The sparsity in (14) can be exploited to derive a sparsity-adapted complex moment-HSOS hierarchy [17]. We solve the second sparse HSOS relaxation of (14) for $l=40,80,\ldots,400$. The results are displayed in Table 5. From the table we see that the reformulation (HSOS-$\mathbb{R}$}) is $1.5\sim 2$ times as fast as the reformulation (HSOS-$\mathbb{R}$}). Moreover, for this problem, solving the original complex SDP with Hypatia is extremely slow probably due to the fact that the SDPs contain many PSD blocks.
@@
$l$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{C}$})
@@ 4.5 Application to the AC-OPF problem
where $V_{i}$ is the voltage, $S_{k}^{g}$ is the power generation, $S_{ij}$ is the power flow (all are complex variables; $\angle\cdot$ stands for the angle of a complex number) and all symbols in boldface are constants. Notice that $G$ is the collection of generators and $N$ is the collection of buses. For a full description on the AC-OPF problem, we refer the reader to [3] as well as [5].
We select test cases from the AC-OPF library PGLiB-OPF [3]. For each case, we solve the minimal relaxation step of the sparse HSOS hierarchy [17]. The results are displayed in Table 6. From the table, we again see that the reformulation (HSOS-$\mathbb{R}$}) is several ($1.4\sim 5$) times as fast as the reformulation (HSOS-$\mathbb{R}$}).
@@
Case $s$ $t$ (HSOS-$\mathbb{R}$}) (HSOS-$\mathbb{R}$})
@@
14_ieee 19 147 2346 $1.9940\text{e}3$ 0.19 422 $1.9940\text{e}3$ 0.10
30_ieee 36 297 4828 $8.1959\text{e}3$ 0.73 836 $8.1960\text{e}3$ 0.37
30_as 36 297 4828 $5.0371\text{e}2$ 0.55 836 $5.0371\text{e}2$ 0.24
39_epri 49 361 5270 $1.3568\text{e}5$ 0.74 966 $1.3579\text{e}5$ 0.54
89_pegase 101 1221 57888 $9.4098\text{e}4$ 63.6 10262 $9.4101\text{e}4$ 15.1
57_ieee 64 563 11102 $3.6644\text{e}4$ 2.36 2008 $3.6644\text{e}4$ 1.06
118_ieee 172 1325 25374 $9.3216\text{e}4$ 8.27 4471 $9.3216\text{e}4$ 2.68
162_ieee_dtc 174 1809 64874 $1.0492\text{e}5$ 43.4 11327 $1.0495\text{e}5$ 13.8
179_goc 208 1827 25712 $6.0859\text{e}5$ 10.3 4368 $6.0860\text{e}5$ 3.57
240_pserc 383 3039 52172 $2.8153\text{e}6$ 31.9 9243 $2.8170\text{e}6$ 10.7
300_ieee 369 2983 53946 $5.3037\text{e}5$ 40.6 9647 $5.3037\text{e}5$ 10.6
500_goc 671 5255 90502 $3.9697\text{e}5$ 89.8 15918 $3.9697\text{e}5$ 25.4
588_sdet 683 5287 79362 $1.9799\text{e}5$ 91.7 13933 $1.9749\text{e}5$ 21.3
793_goc 890 7019 104978 $1.1194\text{e}5$ 105 18536 $1.1222\text{e}5$ 31.5
1888_rte 2178 18257 280580 $1.2537\text{e}6$ 939 47205 $1.2545\text{e}6$ 180
2000_goc 2238 23009 455530 $9.1876\text{e}5$ 2087 77974 $9.1881\text{e}5$ 439
Summary of the benchmark. The benchmark confirms our theoretical findings and demonstrates that our new real reformulation of complex SDPs is indeed more efficient than the classical real reformulation. Especially, the speedup factor gets even bigger as the SDP size grows and may reach several magnitudes on certain instances. It should be also noted that solving the original complex SDP with Hypatia is even slower than solving the classical real reformulation with MOSEK, probably because MOSEK is commercial and the implementations of SDP algorithms based on real numbers are more mature and robust.
@@ Acknowledgments
Funding
This work was jointly funded by National Key R&D Program of China under grant No. 2023YFA1009401 & 2022YFA1005102, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No. 12201618 & 12171324.
Conflict of interest
The authors declare that they have no conflict of interest.
Data availability
The author confirms that all data generated or analysed during this study are included in this article.